Your company has been approached to bid on a contract to sell 5,500 voice recognition (VR) computer keyboards a year for four years. Due to technological improvements, beyond that time they will be outdated and no sales will be possible. The equipment necessary for the production will cost $4.5 million and will be depreciated on a straight-line basis to a zero salvage value. Production will require an investment in net working capital of $465,000 to be returned at the end of the project, and the equipment can be sold for $445,000 at the end of production. Fixed costs are $640,000 per year, and variable costs are $89 per unit. In addition to the contract, you feel your company can sell 14,200, 16,300, 19,300, and 11,800 additional units to companies in other countries over the next four years, respectively, at a price of $198. This price is fixed. The tax rate is 24 percent, and the required return is 10 percent. Additionally, the president of the company will undertake the project only if it has an NPV of $125,000. What bid price should you set for the contract? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

c3 - A B X v fx =SUM(Q2:Q7) C D E F Contributi Variab on per Fixed Sale price le cost unit costs Sales 1 Year units 2 0 3 1 1

08 A B G : x for =SUM(Q2:07) C D E F Contributi Variab on per Fixed Sale price le cost unit costs Depreciati on Sales 1 Year

Hence, Target stock price is $229.54

*Please rate thumbs up